OpenStudy (anonymous):

Integral of (1/(x^3+x))dx?

myininaya (myininaya):

partial fractions! :)

myininaya (myininaya):

\[\frac{1}{x(x^2+1)}=\frac{A}{x}+\frac{Bx+C}{x^2+1}=\frac{A(x^2+1)+x(Bx+C)}{x(x^2+1)}\] => \[1=Ax^2+Bx^2+Cx+A\]

myininaya (myininaya):

\[=>1=A, A+B=0,C=0=>B=-1\]

OpenStudy (ash2326):

\[ \int 1/(x^3+x) dx\] \[ 1/(x^2+1)= 1/x- x/(x^2+1)\]

myininaya (myininaya):

yep :)

OpenStudy (ash2326):

integrate each term \[\int 1/x = ln x\] \[ \int x/(x^2+1)= 1/2 * ln( x^2+1)\]

OpenStudy (anonymous):

Thanks.